Liquid octane will react with gaseous oxygen to produce gaseous carbon dioxide and gaseous water . Suppose 6.85 g of octane is mixed with 9.8 g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Round your answer to 2 significant digits.
1
Expert's answer
2020-06-12T07:26:23-0400

2C8H18+25O2=16CO2+18H2O

1. We can calculate how much CO2 is formed by C8H18 and by O2

  • Based on the above reaction, 114 g of C8H18 to 704 g of CO2 are formed. We calculate how much CO2 is formed from 6.85 g of C8H18 6.85*704/114=42.3gr CO2
  • Based on the above reaction, 800 g of O2 to 704 g of CO2 are formed. We calculate how much CO2 is formed from 9.8 g of O2 9.8*704/800=8.624gr CO2

2.As a final result we get the result calculated on O2 (8.624gr). because when all O2 is consumed, CO2 increases.

if we get the result calculated for CO2 (42.3gr), O2 is not enough for all the CO2 to be consumed. so we do not get this result. and we have the final answer 8.624gr